{2x-y=9
{x-2z=-3
{2y+3z=15
Solve each system!

Answer by stanbon(75887)

{2x-y=9
{x-2z=-3
{2y+3z=15
-------
{2x-y=9
{x-2z= -3
{2y+3z=15
Modify::
y = 2x-9
z = (x+3)/2
------------
Substitute into 2y + 3z = 15 and solve for "x":
2(2x-9) +3(x+3)/2 = 15
------------
4(2x-9) + 3x+9 = 30
8x - 36 + 3x + 9 = 30
11x - 27 = 30
11x = 57
x = 5 2/11
-------
Solve for "y" and for "z"::
y = 2(57/11)-(99/11)
y = 15/11
-------
Solve for "z"::z = (x+3)/2
z = [57/11 + 33/11)/2 = 45/11
===============================
Cheers,
Stan H.
===========